# suppSupplementary References

### v1

@@ Appendix A Adversarial Transferability
In this section, we expand our analysis on adversarial attacks and their transferability across different models. We include more model families, such as DeiT, DenseNet, and VGG. Figures 6, 7, and 8 present the robust accuracy of various models under both white-box and black-box settings for FGSM, PGD, and MIFGSM attacks, respectively. All adversarial examples are crafted with a perturbation budget of $\epsilon=\frac{8}{255}$. For PGD and MIFGSM attacks, we use 20 iterations to craft the adversarial examples.
Appendix B Non Adversarial Analysis
B.1 Drop Along Scanning Direction
@@
Figure 10: Information drop of Tiny and Small family of models along the scanning direction: the image is split into a sequence of fixed-size non-overlapping patches of size 16x16, 8x8. 4x4, and 2x2. We report results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) along the scanning direction. The top row shows results for top-to-bottom (Direction 1) and left-to-right direction(Direction 2). The bottom row shows results for right-to-left (Direction 3) and bottom-to-top direction(Direction 4).
@@
B.2 DINO Drop
In Figure 13 and 14, we report results on salient and non-salient patch drop of information.
B.3 Shuffling Data
B.4 Random Drop
In Table 9 we expand the random patch drop experiments from Table 1 to patch sizes 56$\times$56 and 224$\times$224. Furthermore, in Figure 16, 17, 18, and 21 we expand our analysis on random patch drop to several other models.
@@ B.4 Random Drop
VSSM-T
VSSM-S
VSSM-B
@@
B.5 Results on Domain Generalization
In Table 10 (left) we report results on ImageNet-E dataset across several CNNs, Transformers, and VSSM-based models, and in Table 10 (right), we report results on ImageNet-B dataset. In Table 11 we report results on ImageNetV2 [41], ImageNet-A [19], ImageNet-R [18], and ImageNet-S [11]. These datasets are designed to evaluate the robustness and generalization capabilities of models trained on the original ImageNet dataset. ImageNetV2 [41] tests the models’ performance on previously unseen images that follow the same distribution as the training data, while ImageNet-A [19] contains naturally occurring adversarial examples that are difficult for models to classify correctly. ImageNet-R [18] consists of images with different artistic renditions to test the models’ ability to generalize to different visual domains and styles, and ImageNet-S [11] introduces various types of synthetic corruptions to assess the models’ robustness against image distortions like noise, blur, and compression artifacts.
@@ B.5 Results on Domain Generalization
VSSM-T(v0) 90.03 89.31 70.38 85.59 91.49 89.59 82.01 78.63 95.73 98.00 91.60 94.90 86.10 86.70
VSSM-S(v0) 90.53 90.76 73.39 87.78 93.13 90.71 83.92 80.42 96.16 99.10 92.80 95.40 89.40 88.50
VSSM-B(v0) 91.75 90.62 73.90 88.33 93.29 91.19 83.96 81.66 96.00 98.80 93.50 96.20 89.70 88.40
VSSM-T(v2) 91.15 89.87 75.18 87.41 92.09 91.06 83.66 79.71 95.84 98.50 92.20 96.30 87.20 86.80
VSSM-S(v2) 92.03 90.79 76.15 88.81 93.22 92.25 85.57 81.89 96.37 99.20 94.10 97.40 90.90 89.50
VSSM-B(v2) 92.37 91.27 77.30 89.11 93.70 92.64 86.03 83.62 96.37 99.10 94.00 96.50 90.80 89.80
@@
VSSM-T(v0) 81.92 70.82(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.10)}}}$ 33.01(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.91)}}}$ 13.56(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-68.36)}}}$ 32.11(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-49.81)}}}$ 37.37(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.55)}}}$
VSSM-T(v2) 82.28 71.16(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.12)}}}$ 33.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.29)}}}$ 12.08(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-70.20)}}}$ 32.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-50.23)}}}$ 37.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.96)}}}$
VSSM-S(v0) 83.15 72.89(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.26)}}}$ 38.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.10)}}}$ 17.28(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-65.87)}}}$ 37.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.10)}}}$ 41.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.83)}}}$
VSSM-S(v2) 83.48 73.01(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.47)}}}$ 36.98(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.50)}}}$ 16.45(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-67.03)}}}$ 35.75(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.73)}}}$ 40.55(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.93)}}}$
VSSM-B(v0) 83.48 72.97(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.51)}}}$ 38.24(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.24)}}}$ 18.88(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-64.60)}}}$ 37.33(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.15)}}}$ 41.86(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.62)}}}$
VSSM-B(v2) 83.76 73.22(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.54)}}}$ 38.53(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.23)}}}$ 18.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-65.41)}}}$ 35.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.77)}}}$ 41.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.24)}}}$
B.6 Common Corruptions
In Figure 22 and 23, we present the relative corruption error and mean corruption error (mCE) of various models subjected to all 19 corruption methods applied to the ImageNet dataset. Figure 24 shows the mIoU, mAcc, and aAcc scores on ADE-20K after applying common corruptions to the dataset. Similarly, Figure 25 displays the mAP, APs, APm, and APl scores on the COCO-C dataset. All the scores are averaged across all severity levels of each corruption.
Figure 22: Relative Corruption Error values for different corruptions and architectures on ImageNet-C, with error rates relative to ResNet-50.

### v2

@@ Appendix C Robustness against Adversarial Attacks
In this section, we expand our analysis on adversarial attacks and their transferability across different models. We include more model families, such as DeiT, DenseNet, and VGG. Figures 26, 27, and 28 present the robust accuracy of various models under both white-box and black-box settings for FGSM, PGD, and MIFGSM attacks, respectively. All adversarial examples are crafted with a perturbation budget of $\epsilon=\frac{8}{255}$. For PGD and MIFGSM attacks, we use 20 iterations to craft the adversarial examples.
[…]
@@
Figure 6: Information drop of Tiny and Small family of models along the scanning direction: the image is split into a sequence of fixed-size non-overlapping patches of size 16x16, 8x8. 4x4, and 2x2. We report results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) along the scanning direction. Top row shows results for top-to-bottom (Direction 1) and left-to-right direction(Direction 2). Bottom row shows results for right-to-left (Direction 3) and bottom-to-top direction(Direction 4).
@@
VMamba-T
VMamba-S
VMamba-B
MambaVision-T
MambaVision-S
MambaVision-B
@@
Corruptions $\downarrow$
ResNet-50
ConvNext-T
ConvNext-S
ConvNext-B
ViT-T
ViT-S
ViT-B
VMamba-T
VMamba-S
VMamba-B
MambaVision-T
MambaVision-S
MambaVision-B
Swin-T
Swin-S
Swin-B
Brightness 100.00 68.20 58.63 52.15 223.14 101.09 41.67 61.60 53.06 52.94 64.39 53.73 51.30 75.89 65.35 63.54
Contrast 100.00 40.14 30.29 28.96 426.29 186.49 103.16 38.17 30.02 27.70 52.05 34.93 29.72 44.41 42.09 41.70
Defocus Blur 100.00 85.70 77.82 73.87 103.27 70.64 42.81 84.96 72.32 73.00 77.80 71.71 67.04 87.50 66.82 81.49
Elastic Transform 100.00 81.11 72.18 69.44 70.34 55.69 41.43 82.78 69.89 72.21 78.96 79.82 75.06 86.36 66.53 76.78
Fog 100.00 113.68 86.18 91.19 312.26 119.54 68.79 87.47 64.93 55.88 62.13 46.48 39.35 84.32 87.24 68.61
Frost 100.00 64.12 57.09 50.75 146.53 75.52 27.19 42.05 36.58 30.79 57.65 49.73 40.25 47.77 40.72 38.63
Gaussian Blur 100.00 86.69 84.58 77.75 107.84 82.50 55.22 88.91 77.05 77.21 79.64 74.80 70.72 90.59 81.36 85.37
Gaussian Noise 100.00 52.01 51.36 56.02 130.74 66.75 30.16 56.81 38.29 52.62 115.29 110.36 96.79 73.31 59.86 54.49
Glass Blur 100.00 89.71 85.46 81.43 98.30 84.73 66.52 93.59 86.00 87.28 90.58 90.13 89.29 96.53 86.27 90.24
Impulse Noise 100.00 47.18 44.71 45.78 122.95 61.19 28.27 44.55 31.93 39.72 92.95 84.21 75.73 63.83 48.63 48.68
JPEG Compression 100.00 69.79 59.62 60.96 131.33 78.90 43.49 82.29 65.82 63.30 74.33 69.63 62.72 144.27 88.13 108.22
Motion Blur 100.00 65.05 56.10 51.42 106.50 58.86 33.26 65.84 54.55 54.13 63.42 58.58 53.44 80.88 50.12 62.42
Pixelate 100.00 82.35 71.52 66.00 52.11 27.31 15.88 80.34 66.30 65.02 67.42 67.31 59.34 109.24 100.70 101.51
Saturate 100.00 63.44 53.74 47.61 250.68 116.28 48.06 57.26 47.97 47.07 60.37 52.12 48.11 75.29 62.40 58.34
Shot Noise 100.00 52.69 51.85 56.73 128.00 66.02 30.12 56.20 38.75 51.72 105.73 104.55 91.49 73.51 62.09 57.99
Snow 100.00 54.55 48.57 43.18 134.88 61.12 26.04 54.29 46.02 42.14 54.68 48.88 42.89 57.70 39.38 47.00
Spatter 100.00 39.86 32.09 29.41 117.51 48.16 21.03 33.10 28.59 24.33 39.28 29.69 24.92 32.71 23.79 22.92
Speckle Noise 100.00 44.78 40.91 34.08 132.49 60.64 26.89 42.92 30.43 34.07 84.20 79.68 64.73 61.92 48.85 48.80
Zoom Blur 100.00 82.47 68.22 64.74 125.18 84.89 52.90 82.72 70.04 70.37 76.07 71.43 65.52 93.27 77.34 77.96
mCE 100.00 67.55 59.52 56.92 153.7 79.28 42.26 65.05 53.08 53.76 66.58 50.50 49.61 77.86 63.04 64.98
Figure 18: Corruption Error(CE) values for different corruptions and architectures on ImageNet-C, with error rates relative to ResNet-50.
VMamba-T(v0) 81.92 70.82(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.10)}}}$ 33.01(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.91)}}}$ 13.56(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-68.36)}}}$ 32.11(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-49.81)}}}$ 37.37(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.55)}}}$
VMamba-T(v2) 82.28 71.16(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.12)}}}$ 33.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.29)}}}$ 12.08(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-70.20)}}}$ 32.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-50.23)}}}$ 37.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.96)}}}$
MambaVision-T 82.10 71.50(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.60)}}}$ 33.49(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.61)}}}$ 13.39(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-68.71)}}}$ 32.28(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-49.82)}}}$ 37.66(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.44)}}}$
VMamba-S(v0) 83.15 72.89(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.26)}}}$ 38.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.10)}}}$ 17.28(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-65.87)}}}$ 37.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.10)}}}$ 41.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.83)}}}$
VMamba-S(v2) 83.48 73.01(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.47)}}}$ 36.98(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.50)}}}$ 16.45(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-67.03)}}}$ 35.75(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.73)}}}$ 40.55(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.93)}}}$
MambaVision-S 83.22 72.62(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.60)}}}$ 35.53(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.69)}}}$ 15.97(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-67.25)}}}$ 33.96(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-49.26)}}}$ 39.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-43.70)}}}$
VMamba-B(v0) 83.48 72.97(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.51)}}}$ 38.24(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.24)}}}$ 18.88(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-64.60)}}}$ 37.33(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.15)}}}$ 41.86(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.62)}}}$
VMamba-B(v2) 83.76 73.22(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.54)}}}$ 38.53(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.23)}}}$ 18.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-65.41)}}}$ 35.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.77)}}}$ 41.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.24)}}}$
MambaVision-B 83.96 73.84(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.12)}}}$ 36.69(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.27)}}}$ 21.69(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-62.27)}}}$ 35.72(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.24)}}}$ 41.98(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.98)}}}$
Figure 20: Caliberation Results: Reliability diagrams and ECE on ConvNext-T, ViT-T, Swin-T, VMamba-T, and MambaVision-T across ImageNet, ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A.
Figure 21: Caliberation Results: Reliability diagrams and ECE on ConvNext-S, ViT-S, Swin-S, VMamba-S, and MambaVision-S across ImageNet, ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A.
Figure 22: Caliberation Results: Reliability diagrams and ECE on ViT-B, Swin-B, VMamba-B, and MambaVision-B across ImageNet, ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A.
Figure 23: ECE error across classification models across ImageNet, ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A.
@@
VMamba-T(v0) 90.03 89.31 70.38 85.59 91.49 89.59 82.01 78.63 95.73 98.00 91.60 94.90 86.10 86.70
VMamba-S(v0) 90.53 90.76 73.39 87.78 93.13 90.71 83.92 80.42 96.16 99.10 92.80 95.40 89.40 88.50
VMamba-B(v0) 91.75 90.62 73.90 88.33 93.29 91.19 83.96 81.66 96.00 98.80 93.50 96.20 89.70 88.40
VMamba-T(v2) 91.15 89.87 75.18 87.41 92.09 91.06 83.66 79.71 95.84 98.50 92.20 96.30 87.20 86.80
VMamba-S(v2) 92.03 90.79 76.15 88.81 93.22 92.25 85.57 81.89 96.37 99.20 94.10 97.40 90.90 89.50
VMamba-B(v2) 92.37 91.27 77.30 89.11 93.70 92.64 86.03 83.62 96.37 99.10 94.00 96.50 90.80 89.80
MambaVision-T 90.67 88.83 73.07 86.40 91.93 90.07 81.87 78.42 95.73 98.60 93.70 96.60 89.10 87.70
MambaVision-S 91.22 90.19 75.18 88.19 92.78 91.19 84.01 81.18 96.03 99.40 94.40 97.80 91.40 90.10
MambaVision-B 91.77 90.65 78.42 89.18 93.70 92.81 86.35 83.78 96.30 99.10 94.60 97.20 91.40 90.60
@@
Swin-B 91.08 89.96 75.09 87.87 92.62 91.22 83.43 80.65 95.95 98.60 92.30 95.60 89.20 87.40